# Impact of Data Distribution on Fairness Guarantees in Equitable Deep Learning

### v1

@@ Abstract
We present a comprehensive theoretical framework analyzing the relationship between data distributions and fairness guarantees in equitable deep learning. Our work establishes novel theoretical bounds that explicitly account for data distribution heterogeneity across demographic groups, while introducing a formal analysis framework that minimizes expected loss differences across these groups. We derive comprehensive theoretical bounds for fairness errors and convergence rates, and characterize how distributional differences between groups affect the fundamental trade-off between fairness and accuracy. Through extensive experiments on diverse datasets, including FairVision (ophthalmology), CheXpert (chest X-rays), HAM10000 (dermatology), and FairFace (facial recognition), we validate our theoretical findings and demonstrate that differences in feature distributions across demographic groups significantly impact model fairness, with performance disparities particularly pronounced in racial categories. The theoretical bounds we derive crroborate these empirical observations, providing insights into the fundamental limits of achieving fairness in deep learning models when faced with heterogeneous data distributions. This work advances our understanding of fairness in AI-based diagnosis systems and provides a theoretical foundation for developing more equitable algorithms. The code for analysis is publicly available via https://github.com/Harvard-Ophthalmology-AI-Lab/fairness_guarantees.
@@
Fairness Learning, Equitable Deep Learning, Theoretical Fairness Analysis
@@ 1 Introduction
Recent work has focused on developing fairness-aware learning algorithms and analyzing their theoretical properties. For example, Zietlow et al. [4] formulate the fairness problem as minimizing the absolute difference in the expected per-group error between two groups. Also, the fairness problem in predictive ability has been formalized as minimizing the difference in the expected loss across all demographic groups [5]. This definition provides a mathematical framework for studying fairness in machine learning models and has been used to derive various fairness error bounds and convergence guarantees [6, 7]. However, while existing strategies have improved model fairness to some extent, there has been limited theoretical analysis of the relationship between data distributions and fairness guarantees [8]. The implications of biases are particularly concerning in high-stakes applications, where unfair assessments can perpetuate systemic inequalities, affect individual opportunities, and erode public trust in automated decision systems [9].
@@
Building upon these insights, this work aims to provide a comprehensive theoretical analysis of the impact of data distribution on the fairness guarantees of equitable deep learning models. By leveraging the fairness problem formulation (Definition 3.1) and the theoretical results on fairness error bounds (Theorem 3.4), convergence rates (Theorem 3.14), and group-specific risk bounds (Theorem 3.17), we derive new analytical bounds that explicitly account for the heterogeneity in data distributions across demographic groups. These bounds provide a deeper understanding of the factors that influence the fairness properties of deep learning models and can guide the development of more robust and equitable algorithms.
@@
  • We formalize the fairness problem in terms of minimizing the differences in expected loss across various demographic groups, such as race categories, and introduce assumptions on data distributions and loss functions.
  • We derive a series of theorems that establish fairness error bounds, algorithmic complexity, generalization bounds, convergence rates, and group-specific risk bounds under various assumptions on the data distributions and loss functions. These results shed light on the key factors influencing the fairness properties of the learned models and provide insights into the design of more robust and equitable algorithms.
  • We prove that under certain conditions, the local optima of the fairness problem can outperform those of the supervised learning problem, highlighting the importance of considering fairness criteria in model development.
  • We corroborate our theoretical findings with empirical results on diverse datasets, such as FairVision [11], CheXpert [12] HAM10000 [13] and FairFace [14], demonstrating the impact of data distributions on the equitable performance of deep learning models for tasks such as glaucoma screening.
The rest of this paper is organized as follows. In Section 2, we discuss related work on fairness learning and datasets. Section 3 presents our main theoretical results on the impact of data distribution on fairness guarantees. In Section 4, we discuss the empirical validation of the theoretical findings on FairVision and CheXpert. Section 5 concludes the paper and outlines directions for future research. Additionally, in the Appendix, we provide supplementary theoretical proofs and experimental results on the HAM10000 and FairFace datasets.
@@ 2 Related Work
2.1 Fairness Learning
In recent years, machine learning has achieved exceptional performance across various fields, yet it often produces biased predictions against different demographic groups. To address this issue, fairness learning has been proposed to eliminate or reduce discrimination and bias against certain protected groups, ensuring fair treatment across different groups. Existing fairness learning approaches can generally be categorized into three strategies: pre-processing, which balances the dataset through sampling [15, 16] or generative models [17]; in-processing, which incorporates constraints into the loss function [18]; and post-processing, which transforms the model output to ensure fairness [19, 20].
@@ 2.1 Fairness Learning
2.2 Fairness Dataset
Fairness learning has been extensively studied across various fields, with several fairness datasets being proposed in areas such as finance [23, 24], criminology [25], social sciences [26], and education [27, 28]. Among these fields, healthcare stands out as a critically high-stakes domain where biased predictions can lead to significant and far-reaching consequences. To address fairness comprehensively, we leverage both medical and non-medical datasets to capture a wide range of contexts and challenges.
In the medical domain, several datasets exist for fairness learning [12, 29, 30, 13, 31, 32, 33, 34, 35]. However, most suffer from limitations such as small sample sizes (up to several hundred patients), which hinder robust fairness assessment, or restricted demographic attributes like age and gender, which constrain their utility for comprehensive fairness studies. Among the large-scale datasets particularly suited for fairness learning, the FairVision [11] dataset includes 30,000 fundus photos from 30,000 patients, covering three major eye diseases: age-related macular degeneration, diabetic retinopathy, and glaucoma. It offers extensive demographic attributes, including age, gender, race, ethnicity, preferred language, and marital status, making it highly relevant for fairness studies. The CheXpert [12] dataset provides 224,316 chest radiographs of 65,240 patients with labels for 14 common chest radiographic observations, as well as age, gender, and race attributes essential for fairness learning. Similarly, the HAM10000 dataset includes 9,948 dermatoscopic images of pigmented skin lesions with detailed annotations for seven diagnostic categories and metadata for age and gender, making it suitable for fairness studies in skin lesion classification.
Beyond the medical domain, we also utilize datasets like FairFace, which contains 152,917 facial images from 6,100 unique identities annotated with protected attributes such as gender and skin color. FairFace serves as a critical benchmark for evaluating fairness in computer vision models, enabling a broader exploration of fairness issues in non-medical applications. By incorporating both medical and non-medical datasets, we aim to provide a holistic evaluation of fairness across diverse domains, addressing unique challenges in healthcare while extending the analysis to other high-impact areas.
@@ 3 Main Results
In this section, we present our main theoretical results on the impact of data distribution on the fairness guarantees. We provide a series of theorems that establish fairness error bounds, algorithmic complexity, generalization bounds, convergence rates, and group-specific risk bounds under various assumptions on the data distributions and loss functions. These results shed light on the key factors influencing the fairness properties of the learned models and provide insights into the design of more robust and equitable algorithms.
@@
Theorem 3.2 (Connection with Conventional Fairness).
Let $f$ be a classifier, $\ell$ be a loss function bounded by $M$, and $D_{a_{i}}$ be the data distribution for demographic group $a_{i}$. Let $p_{i}$ be the proportion of positive samples in group $a_{i}$. Then, for any pair of groups $a_{i}$ and $a_{j}$:
$\left|\mathbb{E}_{(x,y)\sim D_{a_{i}}}[\ell(f(x),y)]-\mathbb{E}_{(x,y)\sim D_{a_{j}}}[\ell(f(x),y)]\right|\leq$
$\quad\max\{p_{i},p_{j}\}\cdot\left|L^{+,a_{i}}(f)-L^{+,a_{j}}(f)\right|+2M\left|p_{i}-p_{j}\right|+$
$\quad\min\{1-p_{i},1-p_{j}\}\cdot\left|L^{-,a_{i}}(f)-L^{-,a_{j}}(f)\right|$
where $L^{+,a_{k}}(f)=\mathbb{E}[\ell(f(x),y)|y=1,s=a_{k}]$ is the risk of the positive labeled samples in group $a_{k}$, and $L^{-,a_{k}}(f)=\mathbb{E}[\ell(f(x),y)|y=0,s=a_{k}]$ is the risk of the negative labeled samples in group $a_{k}$. $L^{+,a_{k}}(f)$ and $L^{-,a_{k}}(f)]$ are defined in [36].
@@
First, we decompose the expected loss for each group into positive and negative components:
$\mathbb{E}_{(x,y)\sim D_{a_{k}}}[\ell(f(x),y)]$ $=p_{k}\cdot L^{+,a_{k}}(f)+(1-p_{k})\cdot L^{-,a_{k}}(f)$
where $k\in\{i,j\}$.
Now, consider the difference in expected losses:
@@
$=$ $\quad|(p_{i}\cdot L^{+,a_{i}}(f)+(1-p_{i})\cdot L^{-,a_{i}}(f))-$
$\quad(p_{j}\cdot L^{+,a_{j}}(f)+(1-p_{j})\cdot L^{-,a_{j}}(f))|$
$=$ $\quad|p_{i}\cdot L^{+,a_{i}}(f)-p_{j}\cdot L^{+,a_{j}}(f)+$
$\quad(1-p_{i})\cdot L^{-,a_{i}}(f)-(1-p_{j})\cdot L^{-,a_{j}}(f)|$
$=$ $\quad\left|p_{i}\cdot(L^{+,a_{i}}(f)-L^{+,a_{j}}(f))+(p_{i}-p_{j})\cdot L^{+,a_{j}}(f)\right.$
$\quad\left.+(1-p_{j})\cdot(L^{-,a_{i}}(f)-L^{-,a_{j}}(f))+(p_{j}-p_{i})\cdot L^{-,a_{i}}(f)\right|$
$=$ $\quad\leq p_{i}\cdot\left|L^{+,a_{i}}(f)-L^{+,a_{j}}(f)\right|+|p_{i}-p_{j}|\cdot\left|L^{+,a_{j}}(f)\right|$
$\quad+(1-p_{j})\cdot\left|L^{-,a_{i}}(f)-L^{-,a_{j}}(f)\right|+|p_{j}-p_{i}|\cdot\left|L^{-,a_{i}}(f)\right|$
Note that $p_{i}\leq\max\{p_{i},p_{j}\}$ and $(1-p_{j})\leq\min\{1-p_{i},1-p_{j}\}$. Also, since $\ell$ is bounded by $M$, we have $\left|L^{+,a_{j}}(f)\right|\leq M$ and $\left|L^{-,a_{i}}(f)\right|\leq M$. Apply these bounds:
$\quad\max\{p_{i},p_{j}\}\cdot\left|L^{+,a_{i}}(f)-L^{+,a_{j}}(f)\right|+M|p_{i}-p_{j}|$
$\quad+\min\{1-p_{i},1-p_{j}\}\cdot\left|L^{-,a_{i}}(f)-L^{-,a_{j}}(f)\right|+M|p_{j}-p_{i}|$
$\leq\max\{p_{i},p_{j}\}\cdot\left|L^{+,a_{i}}(f)-L^{+,a_{j}}(f)\right|+2M\left|p_{i}-p_{j}\right|$
$\quad+\min\{1-p_{i},1-p_{j}\}\cdot\left|L^{-,a_{i}}(f)-L^{-,a_{j}}(f)\right|$
@@
Remark. Theorem 3.2 establishes a fundamental connection between the fairness problem 3.1 and conventional group-specific performance metrics, showing that differences in expected loss arise from an interplay between the model’s group-specific performance characteristics and the underlying class distributions. It suggests that achieving fairness requires careful consideration of both these factors, and that fairness assessments should account for natural demographic variations in class distributions.
The positive sample rate (or the proportion of positive samples) plays a significant role in the fairness of machine learning models, particularly in classification tasks. When this rate differs significantly across subgroups, it can introduce or amplify biases, leading to fairness concerns. For example, if a subgroup is underrepresented in the dataset, the model may perform poorly for that group due to insufficient positive examples for training. Conversely, overrepresentation of a subgroup can dominate the model’s learning process, resulting in biased predictions and disparate performance metrics (e.g., accuracy, precision, recall) across groups. Therefore, the positive sample rate is a critical factor in ensuring fairness in machine learning models. Disparities in this rate can lead to biased predictions, unequal outcomes, and reduced trustworthiness of AI systems.
Assumption 3.3 (Bounded Loss and Positive Sample Rate).
Let $r_{i}$ be the positive sample rate for demographic group $a_{i}$, where $\sum_{i=1}^{k}r_{i}=1$. Assume that the loss function $\ell$ is bounded, i.e., $0\leq\ell(\hat{y},y)\leq M$ for some constant $M>0$.
Theorem 3.4 (Fairness Error Bound).
Under Assumption 3.3, given the fairness problem as defined in Definition 3.1, let $f^{*}$ be the optimal function that minimizes the maximum absolute difference in the expected loss across all demographic groups. Let $\hat{f}$ be an estimate of $f^{*}$ based on a finite sample of size n, and assume that $|\hat{f}-f^{*}|_{\infty}\leq\epsilon$ for some $\epsilon>0$. Then, by Hoeffding’s inequality, with probability at least $1-\delta$, the following inequality holds:
$\max_{i,j}|E_{(x,y)\sim D_{a_{i}}}[\ell(\hat{f}(x),y)]-E_{(x,y)\sim D_{a_{j}}}[\ell(\hat{f}(x),y)]|\leq$
$\max_{i,j}|E_{(x,y)\sim D_{a_{i}}}[\ell(f^{(}x),y)]-E_{(x,y)\sim D_{a_{j}}}[\ell(f^{(}x),y)]|+$
$M\sqrt{\frac{\log(2k/\delta)}{2n\min{r_{i}}}}+2L\epsilon$
where $L$ is the Lipschitz constant of the loss function $\ell$.
Proof.
Let $L_{i}(f)=E_{(x,y)\sim D_{a_{i}}}[\ell(f(x),y)]$ be the expected loss for demographic group $a_{i}$ under function $f$. Define $\Delta(f)=\max_{i,j}|L_{i}(f)-L_{j}(f)|$ as the maximum absolute difference in the expected loss across all demographic groups. By the triangle inequality:
$|\Delta(\hat{f})-\Delta(f^{)}|$ $\leq\max_{i,j}|L_{i}(\hat{f})-L_{j}(\hat{f})-(L_{i}(f^{)}-L_{j}(f^{)})|$
$\leq\max_{i,j}(|L_{i}(\hat{f})-L_{i}(f^{)}|+|L_{j}(\hat{f})-L_{j}(f^{)}|)$
For any group $i$, by the Lipschitz continuity of $\ell$:
$|L_{i}(\hat{f})-L_{i}(f^{)}|$ $=|E_{(x,y)\sim D_{a_{i}}}[\ell(\hat{f}(x),y)-\ell(f^{(}x),y)]|$
$\leq E_{(x,y)\sim D_{a_{i}}}[|\ell(\hat{f}(x),y)-\ell(f^{(}x),y)|]$
$\leq LE_{(x,y)\sim D_{a_{i}}}[|\hat{f}(x)-f^{(}x)|]$
$\leq L\epsilon$
Therefore:
$|\Delta(\hat{f})-\Delta(f^{)}|\leq 2L\epsilon$
By Hoeffding’s inequality and the proof from the original Theorem 3.4, with probability at least $1-\delta$:
$\Delta(\hat{f})$ $\leq\Delta(f^{)}+|\Delta(\hat{f})-\Delta(f^{)}|$
$\leq\Delta(f^{)}+M\sqrt{\frac{\log(2k/\delta)}{2n\min{r_{i}}}}+2L\epsilon$
This completes the proof. ∎
Remark. Theorem 3.4 provides an upper bound on the fairness error of the estimated function $\hat{f}$ in terms of the optimal fairness error $\Delta(f^{*})$ and a term that depends on the sample size $n$, the number of demographic groups $k$, the confidence level $\delta$, and the minimum prevalence $\min_{i}r_{i}$. Specifically, the bound suggests that to achieve a smaller fairness error, one should have a larger sample size $n$, a smaller number of demographic groups $k$, a higher confidence level $1-\delta$, and a more balanced distribution of $r_{i}$ across the groups. This result highlights the importance of collecting sufficient and diverse data from each demographic group to ensure equitable performance.

### v2

# On Demographic Group Fairness Guarantees in Deep Learning

@@ Abstract
We present a comprehensive theoretical framework analyzing the relationship between data distributions and fairness guarantees in equitable deep learning. Our work establishes novel theoretical bounds that explicitly account for data distribution heterogeneity across demographic groups, while introducing a formal analysis framework that minimizes expected loss differences across these groups. We derive comprehensive theoretical bounds for fairness errors and convergence rates, and characterize how distributional differences between groups affect the fundamental trade-off between fairness and accuracy. Through extensive experiments on diverse datasets across various modalities (image, tabular data, and text), including FairVision (eye disease detection), CheXpert (pleural effusion detection), HAM10000 (skin lesion classification), FairFace (facial attribute recognition), ACS Income (income prediction), CivilComments-WILDS (toxic comment detection), we validate our theoretical findings and demonstrate that differences in feature distributions across demographic groups significantly impact model fairness, with performance disparities particularly pronounced in racial categories. The theoretical bounds we derive corroborate these empirical observations, providing insights into the fundamental limits of achieving fairness in deep learning models when faced with heterogeneous data distributions. This work advances our understanding of fairness in AI-based diagnosis systems and provides a theoretical foundation for developing more equitable algorithms. Motivated by these theoretical insights, particularly the link between feature distribution shifts and fairness gaps, we propose Fairness-Aware Regularization (FAR), a practical training objective that directly minimizes inter-group discrepancies in feature centroids and covariances to improve equitable performance. We validate the effectiveness of FAR across all datasets considered in this study, consistently observing improvements in overall AUC, ES-AUC, and subgroup performance. The code for analysis is publicly available via https://github.com/Harvard-AI-and-Robotics-Lab/FairnessGuarantee.
@@
Fairness in Machine Learning, Equitable Deep Learning, Theoretical Fairness Analysis
@@ 1 Introduction
Recent work [4, 5, 6, 7] has focused on developing fairness-aware learning algorithms and analyzing their theoretical properties. For example, Zietlow et al. [8] formulate the fairness problem as minimizing the absolute difference in the expected per-group error between two groups. Also, the fairness problem in predictive ability has been formalized as minimizing the difference in the expected loss across all demographic groups [9]. This definition provides a mathematical framework for studying fairness in machine learning models and has been used to derive various fairness error bounds and convergence guarantees [10, 11]. However, while existing strategies have improved model fairness to some extent, there has been limited theoretical analysis of the relationship between data distributions and fairness guarantees [12]. The implications of biases are particularly concerning in high-stakes applications, where unfair assessments can perpetuate systemic inequalities, affect individual opportunities, and erode public trust in automated decision systems [13].
@@
This work establishes a comprehensive theoretical framework linking data distribution heterogeneity across demographic groups to fairness guarantees in deep learning. We derive novel bounds showing that the fairness error—defined as the maximum difference in expected loss across groups—decomposes into irreducible, statistical, and optimization components (Theorem 3.4), and converges at a rate of $O(1/\sqrt{m})$ with sample size $m$ (Theorem 3.14). Crucially, we prove that a group’s excess risk is fundamentally constrained by its feature distribution shift relative to others: specifically, disparities in group-wise means and covariances directly upper-bound performance gaps (Theorems 3.17, 3.19; Corollary 3.20). This formalizes the trade-off between overall accuracy and subgroup equity (Corollary 3.18) and reveals that fairness is limited not just by algorithmic design, but by inherent statistical differences in group data—particularly pronounced in attributes like race. These insights motivate a practical fairness-aware regularization that minimizes inter-group feature discrepancies, thereby tightening the theoretical bounds and improving empirical equity.
@@
  • We formalize fairness as minimizing expected loss discrepancies across demographic groups (e.g., race) and, under assumptions on data distributions and loss functions, derive theoretical guarantees, including fairness error bounds, generalization and group-specific risk bounds, convergence rates, and algorithmic complexity, that elucidate the key drivers of model fairness and inform the design of more robust, equitable algorithms.
  • Motivated by the theoretical insights, we propose Fairness-Aware Regularization (FAR), a practical objective that directly minimizes inter-group discrepancies in feature centroids and covariances, and can be integrated into diverse deep learning models.
  • We validate our theoretical findings with extensive empirical evaluations on six diverse datasets across various modalities (image, tabular data, and text), i.e., FairVision [15], CheXpert [16], HAM10000 [17], FairFace [18], ACS Income [19], and CivilComments-WILDS [20], comprising over 2.5M samples in total. These experiments consistently demonstrate that larger distributional shifts correspond to higher fairness bounds and lower predictive performance, and that our proposed FAR method improves overall AUC, ES-AUC, and subgroup-level equity.
The rest of this paper is organized as follows. In Section 2, we discuss related work on fairness learning and datasets. Section 3 presents our main theoretical results on the impact of data distribution on fairness guarantees. Section 4 introduces our proposed Fairness-Aware Regularization (FAR), derived from the theoretical insights. In Section 5, we discuss the empirical validation of the theoretical findings across all datasets, including FairVision, CheXpert, HAM, FairFace, ACS Income, and CivilComments-WILDS. Section 6 concludes the paper and outlines directions for future research. Table I includes the key symbols and terms used throughout the theoretical sections.
TABLE I: List of Key Notations and Symbols
Symbol Description $x$ Input data (e.g., an image). $y$ True label corresponding to input $x$. $\hat{y}$ Predicted label, $\hat{y}=f(x)$. $a$ Demographic attribute (e.g., race, gender). $a_{i},a_{j}$ Specific values of demographic attribute $a$ (e.g., Asian). $k$ Total number of demographic groups. $f(\cdot)$ Model function mapping input $x$ to prediction $\hat{y}$. $\ell(\hat{y},y)$ Loss function measuring prediction error. $D_{a_{i}}$ Data distribution for demographic group $a_{i}$. $\mathbb{E}_{(x,y)\sim D_{a_{i}}}[\cdot]$ Expectation operator over distribution $D_{a_{i}}$. $\Delta(f)$ Fairness error: $\max_{i,j}\left|\mathbb{E}_{D_{a_{i}}}[\ell(f(x),y)]-\mathbb{E}_{D_{a_{j}}}[\ell(f(x),y)]\right|$. $f^{*}$ Optimal function minimizing fairness error $\Delta(f)$. $\hat{f}$ Estimated/learned function (e.g., empirical risk minimizer). $\epsilon$ Optimality gap for an $\epsilon$-optimal solution. $M$ Upper bound on the loss function: $0\leq\ell(\hat{y},y)\leq M$. $L$ Lipschitz constant of the loss function $\ell$. $\pi_{i}$ Proportion of samples from group $a_{i}$ in the dataset ($\sum_{i=1}^{k}\pi_{i}=1$). $n$ Total sample size. $n_{i}$ Sample size for group $a_{i}$, $n_{i}=n\pi_{i}$. $d_{\text{VC}}(F)$ Vapnik-Chervonenkis (VC) dimension of function class $F$. $R_{i}(f)$ Expected risk for group $a_{i}$: $\mathbb{E}_{(x,y)\sim D_{a_{i}}}[\ell(f(x),y)]$. $R_{\text{emp},i}(f)$ Empirical risk for group $a_{i}$. $R(f)$ Overall fairness risk, often synonymous with $\Delta(f)$. $R_{\text{emp}}(f)$ Empirical fairness risk. $\mu_{i},\Sigma_{i}$ Mean and covariance matrix of the data distribution $D_{a_{i}}$. $\|\cdot\|_{2}$ Euclidean norm. $\|\cdot\|_{F}$ Frobenius norm. $W_{1}(\cdot,\cdot)$ 1-Wasserstein distance between two distributions. $z=f(x)$ Learned feature representation of input $x$. $\bar{z}_{i}$ Centroid (mean) of features for group $a_{i}$: $\mathbb{E}_{(x,y)\sim D_{a_{i}}}[f(x)]$. $\Sigma_{z_{i}}$ Covariance matrix of features for group $a_{i}$. $R_{\text{fair}}(f)$ Proposed fairness regularization term. $\lambda$ Hyperparameter controlling fairness-accuracy trade-off.
@@ 2 Related Work
2.1 Fairness in Learning
In recent years, machine learning has achieved exceptional performance across various fields, yet it often produces biased predictions against different demographic groups. To address this issue, fairness learning [21, 22] has been proposed to eliminate or reduce discrimination and bias against certain protected groups, ensuring fair treatment across different groups [5, 6, 7]. Existing fairness learning approaches can generally be categorized into three strategies: pre-processing, which balances the dataset through sampling [23, 24] or generative models [25]; in-processing, which incorporates constraints into the loss function [26]; and post-processing, which transforms the model output to ensure fairness [27, 28].
@@ 2.1 Fairness in Learning
2.2 Theoretical and Algorithmic Foundations of Fairness
It is generally impossible to satisfy common fairness criteria, such as calibration, balance for the positive class, and balance for the negative class, all at once unless the base rates (i.e., the prevalence of the positive outcome) are identical across demographic groups [31, 12]. Complementing this, [9] proposes to advocate for similarity-based fairness grounded in a task-specific metric, while [27] introduces equality of opportunity, which requires equal true positive rates across protected groups. [28] further explores the calibration-fairness trade-off, showing that enforcing certain fairness constraints can degrade calibration and thereby challenge the reliability of risk scores. These foundational results highlight that fairness interventions must be carefully tailored to domain-specific values and operational goals, as no single definition universally resolves ethical tensions.
Building on these insights, researchers have developed algorithmic frameworks to operationalize fairness constraints [4] and navigate their inherent trade-offs. [11] and [10] propose optimization-based and reduction-based methods, respectively, to enforce group fairness constraints during model training, enabling flexible trade-offs between accuracy and fairness. [32] introduces prejudice remover regularizers, while [33] develops pre-processing techniques to remove disparate impact by transforming features. Causal reasoning has also emerged as a powerful lens: [34] and [35] formalize counterfactual fairness, requiring predictions to remain unchanged under interventions on sensitive attributes, and [36] extends this to path-specific effects, allowing nuanced modeling of direct and indirect discrimination. [37] analyzes the statistical limits of non-discriminatory learning, revealing fundamental sample complexity barriers. More recently, [38] and [39] have urged caution in operationalizing fairness metrics, emphasizing that algorithmic choices embed normative assumptions that must be explicitly justified. In this work, we advance this algorithmic landscape by deriving closed-form theoretical bounds that explicitly link group-specific performance gaps to measurable distributional shifts, specifically the Euclidean distance between group feature centroids and the Frobenius norm of covariance differences. This allows us to propose a novel fairness-aware regularization term that directly minimizes these distributional discrepancies in feature space, thereby tightening the theoretical upper bound on group-wise loss disparity.
@@ 3 Main Results
In this section, we first formalize fairness as minimizing the maximum loss difference across demographic groups (Definition 3.1), then progressively establish how sample size, model complexity, and optimization error affect fairness guarantees (Theorems 3.4, 3.7, 3.10, 3.14). We then show that, under distributional assumptions, the risk for any group is fundamentally bounded by its statistical distance in mean and covariance from other groups (Theorem 3.17). Finally, Theorem 3.19 delivers our core insight: a group’s expected loss is upper-bounded by the population loss plus a term proportional to its feature distribution shift, directly linking measurable data disparities to model unfairness.
@@
Theorem 3.2 (Exact Connection with Conventional Fairness Metrics).
Let $f$ be a classifier, $\ell$ be a loss function, and $D_{a_{i}}$ be the data distribution for demographic group $a_{i}$. Let $p_{i}$ be the proportion of positive samples in group $a_{i}$. Then, for any pair of groups $a_{i}$ and $a_{j}$, the difference in expected loss is exactly given by:
[…]
where $L_{+,a_{k}}(f)=\mathbb{E}[\ell(f(x),y)|y=1,s=a_{k}]$ is the risk of the positive labeled samples in group $a_{k}$, and $L_{-,a_{k}}(f)=\mathbb{E}[\ell(f(x),y)|y=0,s=a_{k}]$ is the risk of the negative labeled samples in group $a_{k}$.
@@
We begin by decomposing the expected loss for each group into contributions from positive and negative samples:
$\mathbb{E}_{(x,y)\sim D_{a_{k}}}[\ell(f(x),y)]=p_{k}\cdot L_{+,a_{k}}(f)+(1-p_{k})\cdot L_{-,a_{k}}(f)$
for $k\in\{i,j\}$. Now, consider the absolute difference in expected losses between groups $a_{i}$ and $a_{j}$:
@@
Substituting the decomposed expressions:
$\begin{split}=&|\left(p_{i}\cdot L_{+,a_{i}}(f)+(1-p_{i})\cdot L_{-,a_{i}}(f)\right)-(p_{j}\cdot L_{+,a_{j}}(f)\\ &+(1-p_{j})\cdot L_{-,a_{j}}(f))|\end{split}$
Rearranging terms:
$|p_{i}\cdot L_{+,a_{i}}(f)-p_{j}\cdot L_{+,a_{j}}(f)+(1-p_{i})\cdot L_{-,a_{i}}(f)-(1-p_{j})\cdot L_{-,a_{j}}(f)|$
This is the desired exact expression, which establishes a direct and precise relationship between the fairness metric (difference in expected loss) and the conventional group-specific performance metrics ($L_{+},L_{-}$) and class distributions ($p_{i},p_{j}$). ∎
Remark. Theorem 3.2 provides an exact decomposition of the fairness metric (the difference in expected loss between groups) into a sum of terms involving group-specific risks ($L_{+,a_{k}}$, $L_{-,a_{k}}$) and class proportions ($p_{k}$). This precise relationship is fundamental because it reveals that fairness disparities are not solely determined by model performance but arise from the interplay of three key factors: (1) differences in positive-sample risk ($L_{+,a_{i}}-L_{+,a_{j}}$), (2) differences in negative-sample risk ($L_{-,a_{i}}-L_{-,a_{j}}$), and (3) differences in the class distribution ($p_{i}-p_{j}$). By stating an equality rather than an inequality, this result avoids the ambiguity associated with loose bounds and arbitrary constants (e.g., $M$), making it a more relevant and precise foundation for analyzing fairness.
Crucially, this decomposition allows Definition 3.1 to be interpreted through the lens of standard decision-based metrics. If $l$ is the 0-1 loss, $\mathbb{E}[l]$ represents the error rate, making Definition 3.1 equivalent to minimizing the Accuracy Gap. Furthermore, since $L_{+,a_{k}}$ and $L_{-,a_{k}}$ correspond to the False Negative Rate ($1-\text{TPR}$) and False Positive Rate (FPR) respectively, Equation (2) demonstrates that minimizing the expected loss difference implicitly bounds the disparities in TPR and FPR. Consequently, our loss-based framework serves as a differentiable surrogate for achieving Equalized Odds and Accuracy Parity in practical deployments.
Assumption 3.3 (Bounded Loss and Group Proportions).
Let $\pi_{i}$ be the proportion of samples for demographic group $a_{i}$, such that in a dataset of size $n$, the number of samples for group $a_{i}$ is $n_{i}=n\pi_{i}$ and $\sum_{i=1}^{k}\pi_{i}=1$. Assume that the loss function $l$ is bounded, i.e., $0\leq l(\hat{y},y)\leq M$ for some constant $M>0$.
Theorem 3.4 (Fairness Error Bound).
Under Assumption 3.3, given the fairness problem as defined in Definition 3.1, let $f^{*}$ be the optimal function that minimizes the maximum absolute difference in the expected loss across all demographic groups, i.e., $f^{*}=\arg\min_{f\in\mathcal{F}}\Delta(f),\quad\text{where}\quad\Delta(f)=\max_{i,j}\left|\mathbb{E}_{D_{a_{i}}}[\ell(f(x),y)]-\mathbb{E}_{D_{a_{j}}}[\ell(f(x),y)]\right|$. Let $\hat{f}$ be an estimate of $f^{*}$ based on a finite sample of size $n$, and assume that $\|\hat{f}-f^{*}\|_{\infty}\leq\epsilon$ for some $\epsilon>0$. Then, with probability at least $1-\delta$, the following inequality holds:
$\max_{i,j}|\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[l(\hat{f}(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{j}}}[l(\hat{f}(x),y)]|\leq$
$\max_{i,j}|\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[l(f^{*}(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{j}}}[l(f^{*}(x),y)]|$
$+M\sqrt{\frac{\log(2k/\delta)}{2n\min_{i}\pi_{i}}}+2L\epsilon$
where $L$ is the Lipschitz constant of the loss function $l$.
Proof.
Let $L_{i}(f)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[l(f(x),y)]$ be the expected loss for demographic group $a_{i}$ under function $f$. Let $\Delta(f)=\max_{i,j}|L_{i}(f)-L_{j}(f)|$ be the maximum absolute difference in the expected loss, which represents the true fairness error.
[…]
$|\Delta(\hat{f})-\Delta(f^{*})|$ $\leq\max_{i,j}|(L_{i}(\hat{f})-L_{j}(\hat{f}))-(L_{i}(f^{*})-L_{j}(f^{*}))|$
$\leq\max_{i,j}(|L_{i}(\hat{f})-L_{i}(f^{*})|+|L_{j}(\hat{f})-L_{j}(f^{*})|)$
For any group $i$, the difference in expected loss is bounded by the Lipschitz property:
$|L_{i}(\hat{f})-L_{i}(f^{*})|$ $=|\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[l(\hat{f}(x),y)-l(f^{*}(x),y)]|$
$\leq\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[|l(\hat{f}(x),y)-l(f^{*}(x),y)|]$
$\leq L\cdot\mathbb{E}_{(x,y)\sim\mathcal{D}_{a_{i}}}[|\hat{f}(x)-f^{*}(x)|]$
Given the assumption that $\|\hat{f}-f^{*}\|_{\infty}\leq\epsilon$, we have:
$|L_{i}(\hat{f})-L_{i}(f^{*})|\leq L\epsilon$
Therefore, the approximation error is bounded by:
$|\Delta(\hat{f})-\Delta(f^{*})|\leq 2L\epsilon$
The bound in the theorem consists of the optimal fairness error $\Delta(f^{*})$, a statistical error term derived from concentration inequalities, and the approximation error term $2L\epsilon$ derived above. By Hoeffding’s inequality, the deviation between empirical estimates and true expectations for all $k$ groups can be bounded. This gives rise to the statistical error term that depends on the sample size of the smallest group, $n\cdot\min_{i}\pi_{i}$.
Combining these components, with probability at least $1-\delta$, we arrive at the final bound:
$\Delta(\hat{f})\leq\Delta(f^{*})+M\sqrt{\frac{\log(2k/\delta)}{2n\min_{i}\pi_{i}}}+2L\epsilon$
@@
Remark. Theorem 3.4 provides a crucial theoretical insight by establishing an upper bound on the fairness error for an estimated model $\hat{f}$. This bound elegantly decomposes the total error into three distinct components: the irreducible fairness error of the best possible model, $\Delta(f^{*})$; a statistical error term, $M\sqrt{\frac{\log(2k/\delta)}{2n\min_{i}\pi_{i}}}$, which quantifies the uncertainty from using a finite sample; and an approximation error, $2L\epsilon$, which accounts for the precision of the optimization algorithm. The statistical error term is particularly instructive, as it shows the bound is fundamentally governed by $\min_{i}\pi_{i}$, the proportion of the smallest demographic group. This highlights that even a massive dataset can yield poor fairness guarantees if one group is severely underrepresented. In essence, the theorem demonstrates that achieving fairness requires a multi-pronged approach: collecting large and representative data from all demographic groups, while also deploying powerful optimization algorithms to find the best possible solution.